OpenStudy (anonymous):

solve: a^5 + 6a^4 + 5a = 0

OpenStudy (anonymous):

a= -1, a = 0

OpenStudy (anonymous):

a^5 + 6a^4 + 5a^3 = 0 First, factor out the a^3 term a^3 * (a^2 + 6a + 5) = 0 Next, we need factors of 5 that add up to 6, easy because 5 is prime... a^3 * (a + 1)(a + 5) = 0 Your values that make this true are a=0, a=-1, a=-5

OpenStudy (anonymous):

does this this help?

OpenStudy (anonymous):

yes...very much thank you

OpenStudy (anonymous):

your welcome:]